# Estimating transformative agreement impact on hybrid open access: A comparative large-scale study using Scopus, Web of Science and open metadata

### v1

@@ 1 Introduction
Transformative agreements are a much-discussed library licensing model for transitioning subscription-based journal publishing to full open access. Although these agreements can vary considerably, they are mainly aimed at hybrid journal bundles, enabling authors from participating institutions to publish open access in these journals while providing reading access to the entire portfolio (Borrego et al., 2021; Hinchliffe, 2019). Following their initial proposal in 2015 (Schimmer et al., 2015), the number of transformative agreements has grown substantially, resulting in increased open access publishing in hybrid journals (Jahn, 2025; McCabe & Mueller-Langer, 2024; Rothfritz et al., 2024). Prior to these agreements, only a few articles in hybrid journals were openly accessible (Björk, 2012; Laakso & Björk, 2016; Piwowar et al., 2018), except for some European countries that had dedicated funding policies and the SCOAP3 consortium for high-energy physics journals (C.-K. (Karl). Huang et al., 2020; Kohls & Mele, 2018; Pinfield, 2015; Robinson-Garcia et al., 2020). Coordination efforts through initiatives such as OA2020 and Efficiency and Standards for Article Charges (ESAC) have further streamlined workflows between library consortia and publishers (Campbell et al., 2022; Geschuhn & Stone, 2017). By April 2025, the ESAC Transformative Agreement Registry, the primary community-driven evidence source, listed approximately 1,300 agreements. [^1] https://esac-initiative.org/about/transformative-agreements/agreement-registry/
@@
These discussions highlight the need for robust data to evaluate transformative agreements. Importantly, these agreements themselves can help improve this evidence base, as researchers and libraries advocate for including open metadata in negotiations to help avoid data analytics from becoming a commodity controlled by publishers (Aspesi & Brand, 2020; McCabe & Mueller-Langer, 2024). For instance, ESAC recommends using Crossref to share metadata supporting workflows including open access licenses and funding information to support discovery and monitoring (Geschuhn & Stone, 2017). Simultaneously, the bibliometric community pushes for more open metadata through this publisher-driven DOI registration platform and continuously monitors progress (van Eck & Waltman, 2024).
This push for comprehensive metadata has yielded results. A growing number of open scholarly data services that build on publisher-provided metadata through Crossref have benefited from increasing metadata coverage, although differences between publishers can be observed (van Eck & Waltman, 2024). OpenAlex stands out as a prominent example that primarily retrieves updated records from Crossref and has gained attention for its inclusivity, openness and comprehensive coverage (Priem et al., 2022). Comparative studies confirm it as a suitable alternative to proprietary bibliometric databases (Alperin et al., 2024; Culbert et al., 2025). Notably, Céspedes et al. (2025) synthesised evidence on both OpenAlex’s strengths and its shortcomings for bibliometric analysis. These include challenges with open access status identification (Jahn et al., 2023; Simard et al., 2024), author attribution (Culbert et al., 2025), institutional affiliations (Zhang et al., 2024), and document type classification (Haupka et al., 2024).
While promising for general bibliometric research, applying open metadata to transformative agreements presents distinct challenges due to their targeted journal coverage and complex funding attribution. With invoicing data largely unavailable, previous studies have relied on publication fee pricing lists or expenditures, and first and/or corresponding author affiliations as proxies to attribute funding to institutions (Butler et al., 2023; Haustein et al., 2024; Zhang et al., 2022). Recent work has extended this approach and attributed open access articles to transformative agreements by using data from the cOAlition S Journal Checker Tool, which links journals and participating institutions to transformative agreements included in the ESAC registry (de Jonge et al., 2025; Jahn, 2025). Yet, no systematic comparison exists of how open versus proprietary data sources perform when analysing the same hybrid journals covered by transformative agreements to identify comparative strengths and shortcomings of data sources.
@@ 2.1 Data sources
hoaddata, developed and maintained by the author to support open access monitoring and research (Achterberg & Jahn, 2023; Jahn, 2025), is an R data package that regularly collects information on hybrid open access uptake from multiple openly available data sources (Jahn, 2024). It combines article-level metadata from Crossref (Hendricks et al., 2020) and affilation metadata from OpenAlex (Priem et al., 2022) with transformative agreement information from the Transformative Agreement Data dump used by the cOAlition S Journal Checker Tool (JCT) [^2] https://journalcheckertool.org/transformative-agreements/, which links journal and institutional data about participating research organisations to agreements in the ESAC Transformative Agreement Registry [^3] https://esac-initiative.org/about/transformative-agreements/agreement-registry/.
@@ 2.2 Data processing steps
Following Jahn (2025), the starting point was a unified dataset of several safeguarded JCT snapshots [^5] https://github.com/njahn82/jct_data. JCT journal data were enriched with ISSN variants linked to ISSN-L according to the ISSN agency. To identify hybrid journals, a comprehensive exclusion of fully open access journals was performed using multiple journal lists including the Directory of Open Access Journals (DOAJ). The resulting hybrid journal data comprising more tahn 13,000 hybrid journals were made available via hoaddata and were used to independently determine the publication volume for each database.
@@
Articles in hybrid journals were considered open access when publishers made them freely available under a CC license on their platforms. While hoaddata obtained this information from Crossref license metadata, Web of Science and Scopus relied on Unpaywall as evidence source. Unpaywall also uses Crossref license metadata, but supplements them by parsing publisher websites directly, addressing cases where publishers do not provide machine-readable CC license information (Piwowar et al., 2018). This additional parsing remains necessary despite transformative agreement workflows recommending the deposition of CC licenses during DOI registration (Geschuhn & Stone, 2017). Both Web of Science [^6] https://webofscience.help.clarivate.com/en-us/Content/open-access.html and Scopus [^7] https://blog.scopus.com/posts/scopus-filters-for-open-access-type-and-green-oa-full-text-access-option defined hybrid open access consistently as content available under CC licenses on publisher platforms, distinguishing it from bronze open access that lack such explicit license information, or use publisher-specific licenses (Piwowar et al., 2018).
@@ 2.3 Data records
As a result of the comprehensive data processing described above, datasets on open access in hybrid journals included in transformative agreements were aggregated for each database at the country and journal levels by year. Table 1 provides a general overview of coverage between 2019 and 2023. It shows that the majority of hybrid journals published at least one original research article or review, collectively marked as “original article” in the following, during the five-year period. These journals have provided the basis for the subsequent calculation of article-level indicators.
@@ 3.1 Coverage comparision
Figure 2 presents the coverage of hybrid journals included in transformative agreements, based on the intersections of the examined data sources. The intersection sets of journals (2A) and articles (2B= are visualised as an UpSet graph (Krassowski, 2020; Lex et al., 2014). The x-axis represents the set intersections in a matrix layout. The analysis included hybrid journals that published at least one open access article between 2019 and 2023, based on open access status information from each database. Only original articles were included in the analysis.
Figure 2: Comparison of hybrid journal indexing by data source, 2019-2023. Only hybrid journals present in the cOAlition S Transformative Agreement Data dump with at least one open access article are considered. A) presents the number of journals, B) the number of articles (DOI), distinguishing between shared DOI corpus and surplus in hoaddata. Box plots (C) shows the five-years journal article volume (log-scale).
@@
Figure 3 presents the coverage of hybrid journals in transformative agreements across data sources from 2019 to 2023 focusing on publisher portfolios. The analysis highlights the dominance of the three largest publishers, Elsevier, Springer Nature, and Wiley, which collectively accounted for 47% of the hybrid journals and 62% of the articles published during this five-year period. In terms of article volume, Elsevier led with 2,441,358 articles (33% of the total) published across 1,951 hybrid journals (16% of the total). Springer Nature followed with 1,247,578 articles (17%) in 2,311 hybrid journals, although it recorded the largest number of hybrid journals (19%). Wiley accounted for 858,939 articles (12%) in 1,382 hybrid journals (11%). The remaining 54 publishers collectively accounted for 2,749,847 articles (38%) in 6,476 hybrid journals (53%).
Figure 3: Comparison of hybrid journal indexing by data source and publisher, 2019-2023. Only hybrid journals present in the cOAlition S Transformative Agreement Data dump with at least one open access article are considered. A) presents the percentage of journals by publisher, B) the number of articles by publisher (shared DOI). Box plots (C) shows the five-years journal article volume (log-scale) by publisher.
@@ 3.2 Open Access Indicator Comparison
Figure 4A shows a moderate growth of open access in hybrid journals, which is consistent across hoaddata (black line), Scopus (yellow line), and Web of Science (blue line). According to hoaddata, hybrid open access uptake increased from 6.4% (n = 85,071) in 2019 to 19% (n = 302,358) in 2023. Similarly, Scopus recorded a growth from 6.5% (n = 84,648) in 2019 to 19% (n = 322,850) in 2023. However, Web of Science recorded higher open access uptake in the early years, before converging to a similar level in 2023, from 11% (n = 137,202) in 2019 to 20% (n = 255,481) in 2023. This suggests a different approach towards labelling hybrid open access by Web of Science.
Similarly, hybrid open access by transformative agreements substantially increased between 2019 and 2023 (Figures 4B and C). Trends were consistent when measuring the first (dashed line) and corresponding authorship (dotted line). According to Scopus, 479,297 open access articles could be attributed to transformative agreements based on first author metadata (increasing from 15,341 to 206,084) and 489,262 using corresponding author metadata (from 15,444 to 210,816). Web of Science recorded 493,028 open access articles via transformative agreements using first author metadata (increasing from 21,871 to 174,126) and 500,076 using corresponding author metadata (from 22,092 to 177,030). hoaddata, lacking corresponding author data, linked 501,649 articles to transformative agreements, increasing from 16,010 in 2019 to 183,757 in 2023.
@@ 4 Discussion
The coverage analysis revealed that hybrid journals are well indexed in all three data sources, particularly in terms of article coverage, which reflects the dominance of major publishers whose established journal portfolios are comprehensively indexed in proprietary databases (Bellen et al., 2024). Differences emerge for journals targeting practitioners or local non-English language communities, with many such titles indexed exclusively in Crossref and Scopus. Using Crossref as a bibliometric database in hoaddata demonstrated a particular strength in identifying newly established hybrid journals, a notable finding given that transformative agreements primarily target existing subscription-based journals. The landscape of hybrid journal publishing thus differs markedly from that of fully open access journals. Comparing the coverage of OpenAlex, Scopus and Web of Science, Simard et al. (2024) indicate that only half of the fully open access journals listed in the DOAJ are also indexed in Scopus and Web of Science. Notably, journals that charge no publication fees (“diamond journals”) are absent from the selective Web of Science, which reinforces existing disparities in the indexing of under-represented research communities and regions in selective bibliometric databases (Simard et al., 2024).
@@
The database comparison revealed important discrepancies in open access evidence. Crucially, not all publishers share CC licence metadata via Crossref, a limitation that becomes apparent when contrasting Crossref with Unpaywall’s data in Scopus and Web of Science. While Unpaywall can detect such gaps by parsing publishers’ websites for open access licences, it nevertheless missed a substantial number of CC-licensed open access articles from Wiley journals indexed in Crossref, likely because of parsing errors on the publisher’s website. This resulted in fewer open access articles being recorded in Scopus and Web of Science. Further inconsistencies emerged between Scopus and Web of Science, despite both relying on Unpaywall: Web of Science erroneously labels Elsevier’s delayed open access as hybrid, whereas Scopus correctly categorises it. Notably, Scopus and Crossref showed greater alignment than that observed in a related comparison between Scopus and OpenAlex (Alperin et al., 2024). This likely reflects the use of a curated list to identify hybrid open access journals, rather than depending solely on article-level tags. Given these variations, which may compromise comparability, research and monitoring exercises should avoid relying on a single source. Instead, the selection of open access evidence should be cross-verified using multiple sources and snapshots that can be used to track changes in the data over time (C.-K. Huang et al., 2020). Incorporating expert-curated journal lists can also help prevent misclassification based on business models (Visser et al., 2021).
@@
It is important to note that the study’s estimates of articles from institutions involved in transformative agreements are approximations because of a lack of access to invoice data, which is not usually shared by library consortia and publishers. Furthermore, the open data sources for transformative agreements, including the cOAlition S Journal Checker Tool and the ESAC registry, are based on voluntary effort, crowd-sourced from various consortia. However, recent validation using Dutch research information demonstrates the reliability of such an open approach for the assessment of articles under transformative agreements (de Jonge et al., 2025). With enhanced open metadata compliance attributable to evolving standards and initiatives in support of negotiations with publishers, particularly through the ESAC initiative and Barcelona Declaration on Open Research Information, the situation is likely to improve. In conclusion, this study has shown that open metadata are well-suited for analysing the transition from subscription-based journal publishing to full open access. But using them in conjunction with proprietary databases provides a more robust understanding that overcomes individual shortcomings in coverage and metadata quality.

### v2

@@ 1 Introduction
Transformative agreements are a much-discussed library licensing model for transitioning subscription-based journal publishing to full open access. Although these agreements can vary considerably, they are mainly aimed at hybrid journal bundles, enabling authors from participating institutions to publish open access in these journals while providing reading access to the entire portfolio (Borrego et al., 2021; Hinchliffe, 2019). Following their initial proposal in 2015 (Schimmer et al., 2015), the number of transformative agreements has grown substantially, resulting in increased open access publishing in hybrid journals (Jahn, 2025; McCabe & Mueller-Langer, 2024; Rothfritz et al., 2024). Prior to these agreements, only a few articles in hybrid journals were openly accessible (Björk, 2012; Laakso & Björk, 2016; Piwowar et al., 2018), except for some European countries that had dedicated funding policies and the SCOAP3 consortium for high-energy physics journals (C.-K. (Karl). Huang et al., 2020; Kohls & Mele, 2018; Pinfield, 2015; Robinson-Garcia et al., 2020). Coordination efforts through initiatives such as OA2020 and Efficiency and Standards for Article Charges (ESAC) have further streamlined workflows between library consortia and publishers (Campbell et al., 2022; Geschuhn & Stone, 2017). By April 2025, the ESAC Transformative Agreement Registry [^1] https://esac-initiative.org/about/transformative-agreements/agreement-registry/, the primary community-driven evidence source, listed around 1,300 agreements between libraries and all the major commercial publishers and leading society publishers(Rothfritz et al., 2024).
@@
These discussions highlight the need for robust data to evaluate transformative agreements. Importantly, these agreements themselves can help improve this evidence base, as researchers and libraries advocate for including open metadata in negotiations to help avoid data analytics from becoming a commodity controlled by publishers (Aspesi & Brand, 2020; McCabe & Mueller-Langer, 2024). For instance, the ESAC Workflow Recommendations for Transformative and Open Access Agreements promote using Crossref to share metadata that supports workflows including open access licenses and funding information to support discovery and monitoring (Geschuhn & Stone, 2017). Simultaneously, the bibliometric community pushes for more open metadata through this publisher-driven DOI registration platform and continuously monitors progress (van Eck & Waltman, 2024).
This push for comprehensive metadata has yielded results. A growing number of open scholarly data services that build on publisher-provided metadata through Crossref have benefited from increasing metadata coverage, although differences between publishers can be observed (van Eck & Waltman, 2024). OpenAlex stands out as a prominent example that primarily retrieves updated records from Crossref and has gained attention for its inclusivity, openness and comprehensive coverage (Priem et al., 2022). Comparative studies confirm it as a suitable alternative to proprietary bibliometric databases (Alperin et al., 2024; Culbert et al., 2025). Notably, Céspedes et al. (2025) synthesised evidence on both OpenAlex’s strengths and its shortcomings for bibliometric analysis. These include challenges with open access status identification (Jahn et al., 2023; Simard et al., 2025), author attribution (Culbert et al., 2025), institutional affiliations (Zhang et al., 2024), and document type classification (Haupka et al., 2024).
While promising for general bibliometric research, measuring hybrid open accesss and transformative agreements presents particular challenges. With invoicing data largely unavailable (Kramer, 2024), previous studies have relied on publication fee pricing lists or expenditures, and first and/or corresponding author affiliations as proxies to attribute funding to institutions (Butler et al., 2023; Haustein et al., 2024; Zhang et al., 2022). Others manually obtained journals and institutions from transformative agreement documents before retrieving eligible publications from the Web of Science (Bakker et al., 2024). Recent work has extended this approach and programmatically attributed open access articles to transformative agreements by using data from the cOAlition S Journal Checker Tool, which links journals and participating institutions to transformative agreements included in the ESAC registry, and the open metadata sources Crossref and OpenAlex (de Jonge et al., 2025; Jahn, 2025). Yet, no systematic comparison exists of how open versus proprietary bibliometric data sources perform when analysing the same hybrid journals covered by transformative agreements to identify comparative strengths and shortcomings of data sources.
@@ 2.1 Data sources
hoaddata [^2] https://github.com/subugoe/hoaddata, developed and maintained by the author to support open access monitoring and research (Achterberg & Jahn, 2023; Jahn, 2025), is an R data package that regularly collects information on hybrid open access uptake from multiple openly available data sources (Jahn, 2024). It combines article-level metadata from Crossref (Hendricks et al., 2020) and affilation metadata from OpenAlex (Priem et al., 2022) with transformative agreement information from the Transformative Agreement Data dump used by the cOAlition S Journal Checker Tool (JCT) [^3] https://journalcheckertool.org/transformative-agreements/, which links journal and institutional data about participating research organisations to agreements in the ESAC Transformative Agreement Registry [^4] https://esac-initiative.org/about/transformative-agreements/agreement-registry/.
@@ 2.2 Data processing steps
Following Jahn (2025), the starting point was a unified dataset of several safeguarded JCT snapshots [^6] https://github.com/njahn82/jct_data. JCT journal data were enriched with ISSN variants linked to ISSN-L according to the ISSN agency. To identify hybrid journals, a comprehensive exclusion of fully open access journals was performed using multiple journal lists including the Directory of Open Access Journals (DOAJ). The resulting hybrid journal data comprising more than 13,000 hybrid journals were made available via hoaddata and were used to independently determine the publication volume for each database.
@@
Articles in hybrid journals were considered open access when publishers made them freely available under a CC license on their platforms. While hoaddata obtained this information from Crossref license metadata, Web of Science and Scopus relied on Unpaywall as evidence source. Unpaywall also uses Crossref license metadata, but supplements them by parsing publisher websites directly, addressing cases where publishers do not provide machine-readable CC license information (Piwowar et al., 2018). This additional parsing remains necessary despite transformative agreement workflows recommending the deposition of CC licenses during DOI registration (Geschuhn & Stone, 2017). Both Web of Science [^7] https://webofscience.help.clarivate.com/en-us/Content/open-access.html and Scopus [^8] https://web.archive.org/web/20250426153752/https://blog.scopus.com/posts/scopus-filters-for-open-access-type-and-green-oa-full-text-access-option defined hybrid open access consistently as content available under CC licenses on publisher platforms, distinguishing it from bronze open access that lack such explicit license information, or use publisher-specific licenses (Piwowar et al., 2018).
@@ 2.3 Data records
As a result of the comprehensive data processing described above, datasets on open access in hybrid journals included in transformative agreements were aggregated for each database at the country and journal levels by year. Table 1 provides a general overview of coverage between 2019 and 2023. The table categorises hybrid journal coverage in terms of publishing activity. Journals with at least one publication during the five-years period (“active journals”) were further differentiated based on whether they published at least one original research article or review (“active journals (original)”), collectively marked as “original article” in the following. The table also presents the number of hybrid journals that made at least one original article open access (“active journals (original) with OA”). These journals formed the basis for subsequent calculations of article-level indicators.
@@ 3.1 Coverage comparision
Figure 2 presents the coverage of hybrid journals included in transformative agreements, based on the intersections of the examined data sources. The intersection sets of journals (2A) and articles (2B) are visualised as an UpSet graph (Krassowski, 2020; Lex et al., 2014). The x-axis represents the set intersections in a matrix layout. The analysis included hybrid journals that published at least one open access article between 2019 and 2023, based on open access status information from each database. Only original articles were included in the analysis.
Figure 2: Comparison of hybrid journal indexing by data source, 2019-2023. Only hybrid journals present in the cOAlition S Transformative Agreement Data dump with at least one open access article are considered. A) presents the number of journals, B) the number of articles (DOI), distinguishing between shared DOI corpus and surplus in hoaddata. Box plots (C) shows the five-years journal article volume (log scale). Note that intersections sets with at least 30 journals are shown.
@@
Figure 3 presents the coverage of hybrid journals in transformative agreements across data sources from 2019 to 2023 with a focus on publisher portfolios. The analysis highlights the dominance of the three largest publishers, Elsevier, Springer Nature, and Wiley, which collectively accounted for 47% of the hybrid journals and 62% of the articles published during this five-year period. In terms of article volume, Elsevier led with 2,441,358 articles (33% of the total) published across 1,951 hybrid journals (16% of the total). Springer Nature followed with 1,247,578 articles (17%) in 2,311 hybrid journals, although it recorded the largest number of hybrid journals (19%). Wiley accounted for 858,939 articles (12%) in 1,382 hybrid journals (11%). The remaining 54 publishers collectively accounted for 2,749,847 articles (38%) in 6,476 hybrid journals (53%).
Figure 3: Comparison of hybrid journal indexing by data source and publisher, 2019-2023. Only hybrid journals present in the cOAlition S Transformative Agreement Data dump with at least one open access article are considered. A) presents the percentage of journals by publisher, B) the number of articles by publisher (shared DOI). Box plots (C) shows the five-years journal article volume (log scale) by publisher. Note that intersections sets with at least 30 journals are shown.
@@ 3.2 Open Access Indicator Comparison
Figure 4A shows a moderate growth of open access in hybrid journals, which is consistent across hoaddata (black line), Scopus (yellow line), and Web of Science (blue line). According to hoaddata, hybrid open access uptake increased from 6.4% (n = 85,071) in 2019 to 19% (n = 302,358) in 2023. Similarly, Scopus recorded a growth from 6.5% (n = 84,648) in 2019 to 19% (n = 322,850) in 2023. However, Web of Science recorded higher open access uptake in the early years, before converging to a similar level in 2023, from 11% (n = 137,202) in 2019 to 20% (n = 255,481) in 2023. This suggests that Web of Science takes a different approach to labelling hybrid open access, which will become clearer when results by publisher are presented later.
Figures 4B and C highlight that hybrid open access by transformative agreements substantially increased between 2019 and 2023. Trends were consistent when measuring first (dashed line) and corresponding authorship (dotted line). According to Scopus, 479,297 open access articles could be attributed to transformative agreements based on first author metadata (increasing from 15,341 to 206,084) and 489,262 using corresponding author metadata (from 15,444 to 210,816). Web of Science recorded 493,028 open access articles via transformative agreements using first author metadata (increasing from 21,871 to 174,126) and 500,076 using corresponding author metadata (from 22,092 to 177,030). hoaddata, lacking corresponding author data, linked 501,649 articles to transformative agreements, increasing from 16,010 in 2019 to 183,757 in 2023.
@@ 4 Discussion
The coverage analysis revealed that hybrid journals are well indexed in all three data sources, particularly in terms of article coverage, which reflects the dominance of major publishers whose established journal portfolios are comprehensively indexed in proprietary databases (Bakker et al., 2024; van Bellen et al., 2025). Differences emerge for journals targeting practitioners or local non-English language communities, with many such titles indexed exclusively in Crossref and Scopus. Using Crossref as a bibliometric database in hoaddata demonstrated a particular strength in identifying newly established hybrid journals, a notable finding given that transformative agreements primarily target existing subscription-based journals. The landscape of hybrid journal publishing thus differs markedly from that of fully open access journals. Comparing the coverage of OpenAlex, Scopus and Web of Science, Simard et al. (2025) indicate that only half of the fully open access journals listed in the DOAJ are also indexed in Scopus and Web of Science. Notably, journals that charge no publication fees (“diamond journals”) are absent from the selective Web of Science, which reinforces existing disparities in the indexing of under-represented research communities and regions in selective bibliometric databases (Simard et al., 2025).
@@
The database comparison revealed important discrepancies in open access evidence. Crucially, not all publishers share CC licence metadata via Crossref, a limitation that becomes apparent when contrasting Crossref with Unpaywall’s data in Scopus and Web of Science. While Unpaywall can detect such gaps by parsing publishers’ websites for open access licences, it nevertheless missed a substantial number of CC-licensed open access articles from Wiley journals indexed in Crossref, likely because of parsing errors on the publisher’s website. This resulted in fewer open access articles being recorded in Scopus and Web of Science. Further inconsistencies emerged between Scopus and Web of Science, despite both relying on Unpaywall: Web of Science erroneously labels Elsevier’s delayed open access as hybrid, whereas Scopus correctly categorises it. Notably, Scopus and hoaddata showed greater alignment than that observed in a related comparison between Scopus and OpenAlex (Alperin et al., 2024). This likely reflects the advantage of using of a curated list to identify hybrid journals, rather than depending solely on article-level open access tags.
The observed variations may compromise comparability not just in research, but also across open access monitoring services, which have grown over the years (Pampel & Schneider, 2025; Salamoura & Tsakonas, 2024). This study highlights, in line with previous research (Salamoura & Tsakonas, 2024), that design decisions underlying these services, such as data sources used, update frequencies, and open access definitions, are critical. Stakeholders should therefore advocate for greater transparency from monitoring service providers in their methodological choices and data processing decisions (Kramer, 2024). Research and monitoring exercises might also benefit from avoiding reliance on a single source. Instead, the selection of open access evidence could be cross-verified using multiple sources and snapshots that can be used to track changes in the data over time (C.-K. Huang et al., 2020; Pölönen et al., 2020). Incorporating expert-curated journal lists may also help reduce misclassification based on business models (Visser et al., 2021).
@@
It should be noted that the study’s estimates of articles from institutions involved in transformative agreements are approximations because of a lack of access to invoice data, which is not usually shared by library consortia and publishers (Kramer, 2024). Furthermore, the open data sources for transformative agreements, including the cOAlition S Journal Checker Tool and the ESAC registry, are based on voluntary effort, crowd-sourced from various consortia. Additionally, while this study observed strong correlations at the country level, institutional-level correlations may differ due to matching inaccuracies that can occur when working with long lists of participating institutions and different disambiguation approaches for author affiliations. Future research could explore institutional-level analysis in more detail, though recent validation using Dutch research information demonstrates the reliability of such an open approach for assessing articles under transformative agreements (de Jonge et al., 2025).
Despite these current limitations, the situation is likely to improve with enhanced open metadata compliance attributable to evolving standards and initiatives in support of negotiations with publishers, particularly through the ESAC initiative and Barcelona Declaration on Open Research Information, the situation is likely to improve. For example, AIP Publishing and Emerald, whose hybrid journal portfolios could not be included in this study due to a lack of CC licence information, have recently begun sharing this metadata via Crossref. Furthermore, during the 17th Berlin Open Access Conference (B17), an international conference organised by the OA2020 Initiative, research organisations and their library consortia emphasised the global importance of open research information for achieving transparency. In conclusion, this study has shown that open metadata are well-suited for analysing the transition from subscription-based journal publishing to full open access. But using them in conjunction with proprietary databases provides a more robust understanding that overcomes individual shortcomings in coverage and metadata quality.
Acknowledgement
I would like to thank the two reviewers and the PREreview reviewers (Rittman et al., 2025) for their valuable and helpful feedback on the manuscript.